OpenStudy (anonymous):

For f(x) = x3 – 3x2, determine the number and type of possible complex zeros, and find the zeros to the nearest tenth.

OpenStudy (anonymous):

two real roots; 6, 0 three real roots; 3, 0, 0 three real roots; 3, 1, 2 four real roots; 3, 0, 1, –1

OpenStudy (amorfide):

let f(x)=0 x^3 - 3x²=0 take out the highest factor x²(x-3)=0 can you do it now?

OpenStudy (anonymous):

I don' tknow how to find the root from that. :\

OpenStudy (amorfide):

you have two equations x²=0 x-3=0 work out x in both cases

OpenStudy (anonymous):

oh. okay. i get it. thanks.